# Errors-In-Variables Model Fitting for Partially Unpaired Data Utilizing Mixture Models

### v1

@@ Abstract
The goal of this paper is to introduce a general argumentation framework for regression in the errors-in-variables regime, allowing for full flexibility about the dimensionality of the data, error probability density types, the (linear or nonlinear) model type and the avoidance of explicit definition of loss functions. Further, we introduce in this framework model fitting for partially unpaired data, i.e. for given data groups the pairing information of input and output is lost (semi-supervised). This is achieved by constructing mixture model densities, which directly model this loss of pairing information allowing for inference. In a numerical simulation study linear and nonlinear model fits are illustrated as well as a real data study is presented based on life expectancy data from the world bank utilizing a multiple linear regression model. These results allow the conclusion that high quality model fitting is possible with partially unpaired data, which opens the possibility for new applications with unfortunate or deliberate loss of pairing information in the data.
@@ 1 Introduction
Parametric model fitting is a standard task in many applications starting from problem specific models with a few meaningful parameters to huge, flexible models (such as in the training phase of ANNs) [1, 2]. The general idea is that a defined parametric model family is fitted to given input / output data (i.e. supervised learning) with the purpose to estimate the best fit parameters of the model representing the data. Typically, loss functions are defined for this purpose, such as the squared or absolute losses [3]. Practical difficulties of model fitting are i) finding the appropriate model family and their parameters, ii) defining an adequate loss function, iii) applying an efficient optimization algorithm and iv) dealing with data deficiencies (such as outliers, strong noise, incompleteness, partially lost pairing information, etc.).
@@
A possibility to avoid the definition of an ad-hoc loss function is the modeling of known uncertainties in the data and applying Maximum Likelihood (ML) approaches, which, in consequence, lead inherently to data driven loss functions. A well-known connection is between a normal distributed error in the output and the squared loss function. These approaches can always be extended to Bayesian estimations if prior distributions are assumed for the model parameters leading to Maximum A Posteriori (MAP) or Minimum Mean Squared Error (MMSE) algorithms based on the posterior distribution [4].
@@
Unpaired / unlabeled data (also broken sampling in regression) can occur in many applications and is focus of current research, e.g. [5, 6, 7]. In this work, we understand by partially unpaired data (under the term semi-supervised data) specifically that for parts of the data the one-to-one pairing of input and output data is missing, but we still have paired subgroups in the data. Semi-supervised data often is understood as having additional input data without output data, e.g. [8, 9]. This is different to the partially unpaired data in this paper and we regard this as a different flavor of semi-supervised data since we lose significant information compared to a fully supervised framework, but we still have (possibly weak) input/output relations in contrast to unsupervised learning. See Figure 1 for an illustration. It is demonstrated that mixtures of labeled and unlabeled data can improve the predictive performance in regression problems [6]. There are different strategies to deal with such deficiencies, such as altering the data set by data imputation approaches, e.g. see [10, 11], or focusing on unordered subsets in hypothesis testing [7]. We regard the work of Liang et al. [6] considering a general predictive Bayesian frameworks for mixed labeled and unlabeled data as closest to our goal. Although the work provides a very general framework for regression and classification, it misses the configurational complexity of partially unpaired data. In consequence, one task of this paper is to incorporate the largest possible variety of missing pairing information transparently to model fitting by avoiding data alteration (deletion or imputation) and actively constructing mixture model probability densities accurately representing the partial pairing.
@@
Mixture Models are applied in a variety of applications, especially Gaussian Mixture Models (GMM) are very popular, e.g. [6]. A main task is in the literature to find the GMM description of a given data set utilizing expectation maximization (EM) algorithms, also for the task of missing data deficiencies [12, 13]. To the knowledge of the authors, a general framework for constructing mixture models for dealing with the high complexity of corrupt / incomplete pairing information of the data with the goal to support a model fitting problem is missing in the literature. Closely related approaches of utilizing mixture models in applied mathematics with lost pairing information are presented for computer vision in order to deal with unknown feature correspondences [14] and for random equations with high combinatorial possibilities for the stochastic parameters [15].
@@
In this paper, we propose to consider the problem of model fitting in a new conclusive way. A general probabilistic framework for fitting models in data based purely on observational probability density functions including errors-in-variables is presented, which incorporates well-known standard methods for completely paired data, such as ordinary least squares, total least squares (e.g. Deming regression [16]), interval data regression as well as multiple linear regression. This framework will be generalized from supervised to semi-supervised model fitting by including (partially) unpaired data in one common line of stochastic argumentation. The derivations are allow for flexibility about i) the number and dimensions of the input / output data, ii) the type of individual error characteristics of each data point with errors-in-variables utilizing general density functions, iii) the type of (linear or nonlinear) models which should be fitted and iv) the pairing information level. In summary, there are two main goals of this paper: A) presenting a general stochastic argumentation framework for model fitting (without ad-hoc loss functions) and B) presenting an extension of model fitting within this framework to partially unpaired data utilizing mixture models.
@@
  • completely paired data sets as the standard case in model fitting (i.e. supervised learning) (Section 2.2).
@@
  • partially unpaired data sets, which lie between the two previous extremes and include different levels of pairing (i.e. semi-supervised learning) (Section 2.4).
@@ 2.2 Model Fit with Completely Paired Data
We introduce disturbances in the input and output data with the notation (which is referred in literature to errors-in-variables [17]):
@@
2.2.1 Example: Fitting a Line and Gaussian Disturbance (Deming Regression)
In the line of total least squares [17], we are introducing Gaussian disturbances by ${\eta}_{l}\sim\mathcal{N}(0,\,\sigma_{{\eta}}^{2})(s)\;\forall l=1,..,L$ and ${\varepsilon}_{l}\sim\mathcal{N}(0,\,\sigma_{{\varepsilon}}^{2})(s)\;\forall l=1,..,L$. Utilizing the one-dimensional affine model $M(x;\boldsymbol{\alpha})=\alpha_{1}+\alpha_{2}\cdot x$ and inserting it in Equation (4), we get
@@ 2.2.1 Example: Fitting a Line and Gaussian Disturbance (Deming Regression)
=$\quad$= $\text{argmax}_{\;\boldsymbol{\alpha}}\quad\prod\limits_{l=1}^{L}\;\text{e}^{-\frac{(\alpha_{1}+\alpha_{2}\cdot x_{l}-y_{l})^{2}}{2\,(\alpha_{2}^{2}\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2})}}$
=$\quad$= $\text{argmin}_{\;\boldsymbol{\alpha}}\quad\sum\limits_{l=1}^{L}\;\frac{(\alpha_{1}+\alpha_{2}\cdot x_{l}-y_{l})^{2}}{2\,(\alpha_{2}^{2}\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2})}\;.$
The solution to this minimization is the classical unique Deming regression, also known as orthogonal or total least squares, as presented in Supplement B. This argment can be extended to fitting hyperplanes, i.e. errors-in-variables multiple linear regression, as presented also in Supplement B.
@@ 2.2.2 Connection to Interval Data Regression
Interval data is defined as data for which only the borders of an interval in which the true data point lies are observed. Performing model fitting for such data is an active field of research, e.g. for multilinear linear regression models [18, 19].
@@ 2.3 Model Fit with Completely Unpaired Data
In this section, we are losing the property of tuples, i.e. observations of type $\boldsymbol{x}_{h}$ for $h=1,..,H$ and $\boldsymbol{y}_{l}$ for $l=1,..,L$ are unpaired. Only a set of $\boldsymbol{x}_{h}$ and a set of $\boldsymbol{y}_{l}$ observations are available. Please note, even the sizes $H$ and $L$ can be different.
Although there is theoretical research about the usability of such broken sampling data sets, e.g. [5], obviously such data will only lead to very limited regression results if there are no further assumptions about the involved probability densities since we only have marginal distributions. Nonetheless, we want to introduce a formulation by mixture models for this case, which will later be utilized for partially unpaired data directly.
We start this argumentation with possible disturbances in input and output data:
@@
Remark: This fit with completely unpaired data is practically useless. This means, there will be broad, probably non-distinct or multiple maxima in this objective function. Still, this argumentation helps in a theoretical perspective since it is applied directly to the partially unpaired data where we have a range of different levels of pairing information. In Supplement C concrete examples for completely unpaired data model fits are presented mathematically.
@@ 2.4 Model Fit with Partially Unpaired Data
This section introduces the argumentation for partially unpaired data, which is a main result of this paper. For this case, we partition the $H$ observations $\boldsymbol{x}_{h}$ and $L$ observations $\boldsymbol{y}_{l}$ into $r=1,..,R$ disjoint independent groups, i.e. observations of group $r$ of type $\boldsymbol{x}_{h,r}$ for $h=1,..,H_{r}$ and $\boldsymbol{y}_{l,r}$ for $l=1,..,L_{r}$ are unpaired with $H_{r}$ and $L_{r}$ representing the number of elements in subgroup $r$. This means, we have a set of $\boldsymbol{x}_{h}$-values and a set of $\boldsymbol{y}_{l}$-values for each subgroup and we have pairing information on the group level. The number of input and output elements in each subgroup $H_{r}$ and $L_{r}$ are not necessarily the same.
Again, we allow disturbances in $\boldsymbol{x}_{h}$ and $\boldsymbol{y}_{l}$:
@@
Remark: In the partially unpaired setup, we always work with input / output correspondences, only on a subgroup basis. A completely paired data subset (= supervised data) is represented by subgroups of size one. An additional pure input data subset (= unsupervised) can be approximated by neglecting the information about $\boldsymbol{Y}^{\ast}_{r}$, which corresponds to extremely flat $f_{\boldsymbol{\varepsilon}_{l,r}}$, arriving at the classical definition of semi-supervised. This means, in Equation (9) the first term in the integral $f_{\boldsymbol{\varepsilon}_{l,r}}$ gets essentially constant (independent of the prediction $\boldsymbol{M}(\boldsymbol{s};\boldsymbol{\alpha})$) and, therefore, this part becomes practically noninformative with respect to the optimization on $\boldsymbol{\alpha}$. This means, as expected, the completely unsupervised part of the data only on the input side can be neglected since it contains no information about the model parameters $\boldsymbol{\alpha}$.
Remark: An interesting point is that the same cannot be said about having unsupervised data on the output side, e.g. neglecting information about $\boldsymbol{X}^{\ast}_{r}$. This time $f_{\boldsymbol{\eta}_{h,r}}$ becomes a flat distribution, essentially leaving $f_{\boldsymbol{\varepsilon}_{l,r}}(\boldsymbol{y}_{l,r}-\boldsymbol{M}(\boldsymbol{s};\boldsymbol{\alpha}))$ in the integral of Equation (9) evaluated for all possible $\boldsymbol{s}$. This time the change of $\boldsymbol{\alpha}$ can have direct influence on the optimization, essentially taking care that the output values $\boldsymbol{y}_{l,r}$ are plausible / possible (i.e. in the probabilistically blurred image of $\boldsymbol{M}(\boldsymbol{s};\boldsymbol{\alpha})$) for a given parameter set $\boldsymbol{\alpha}$. This shows an asymmetry with respect the classical semi-supervised setup [6].
2.4.1 Example: Fitting a Line and Gaussian Disturbance (Deming Regression)

### v2

@@ Abstract
The goal of this paper is to introduce a general argumentation framework for regression in the errors-in-variables regime, allowing for full flexibility about the dimensionality of the data, error probability density types, the (linear or nonlinear) model type and the avoidance of ad-hoc definitions of loss functions. In this framework we introduce model fitting for partially unpaired data, i.e. for given data groups the pairing information of input and output is lost (semi-supervised). This is achieved by constructing mixture model densities, which directly model this loss of pairing information allowing for inference. In a numerical simulation study linear and nonlinear model fits are illustrated as well as a real data study is presented based on life expectancy data from the world bank utilizing a multiple linear regression model. These results allow the conclusion that high quality model fitting is possible with partially unpaired data, which opens the possibility for new applications with unfortunate or deliberate loss of pairing information in the data.
@@ 1 Introduction
Parametric model fitting is a standard task in many applications starting from problem specific models with a few meaningful parameters to huge, flexible models (such as in the training phase of ANNs) [Zhang, 1997, Bishop, 2006]. The general idea is that a defined parametric model family is fitted to given input / output data (i.e. supervised learning) with the purpose to estimate the best fit parameters of the model representing the data. Typically, loss functions are defined for this purpose, such as the squared or absolute losses [Wang et al., 2022a]. Practical difficulties of model fitting are i) finding the appropriate model family and their parameters, ii) defining an adequate loss function, iii) applying an efficient optimization algorithm and iv) dealing with data deficiencies (such as outliers, strong noise, incompleteness, partially lost pairing information, etc.).
@@
A possibility to avoid the definition of an ad-hoc loss function is the modeling of known uncertainties in the data and applying Maximum Likelihood (ML) approaches, which, in consequence, lead inherently to data driven loss functions. A well-known connection is between a normal distributed error in the output and the squared loss function. These approaches can always be extended to Bayesian estimations if prior distributions are assumed for the model parameters leading to Maximum A Posteriori (MAP) or Minimum Mean Squared Error (MMSE) algorithms based on the posterior distribution [Hoegele et al., 2013].
@@
Unpaired / unlabeled data (also broken sampling in regression) can occur in many applications and is focus of current research, e.g. [Bai and Hsing, 2005, Liang et al., 2007, Wang et al., 2022b]. In this work, we understand by partially unpaired data (under the term semi-supervised data) specifically that for parts of the data the one-to-one pairing of input and output data is missing, but we still have paired subgroups in the data. Semi-supervised data often is understood as having additional input data without output data, e.g. [Kostopoulos et al., 2018, Qi and Luo, 2022]. This is different to the partially unpaired data in this paper and we regard this as a different flavor of semi-supervised data since we lose significant information compared to a fully supervised framework, but we still have (possibly weak) input/output relations in contrast to unsupervised learning. See Figure 1 for an illustration. It is demonstrated that mixtures of labeled and unlabeled data can improve the predictive performance in regression problems [Liang et al., 2007]. There are different strategies to deal with such deficiencies, such as altering the data set by data imputation approaches, e.g. see [Bennett, 2001, Sterne et al., 2009], or focusing on unordered subsets in hypothesis testing [Wang et al., 2022b]. We regard the work of Liang et al. [Liang et al., 2007] considering a general predictive Bayesian frameworks for mixed labeled and unlabeled data as closest to our goal. Although the work provides a very general framework for regression and classification, it misses the configurational complexity of partially unpaired data. In consequence, one task of this paper is to incorporate the largest possible variety of missing pairing information transparently to model fitting by avoiding data alteration (deletion or imputation) and actively constructing mixture model probability densities accurately representing the partial pairing.
@@
Mixture Models are applied in a variety of applications, especially Gaussian Mixture Models (GMM) are very popular, e.g. [Liang et al., 2007]. A main task is in the literature to find the GMM description of a given data set utilizing expectation maximization (EM) algorithms, also for the task of missing data deficiencies [Michael et al., 2020, McCaw et al., 2022]. To the knowledge of the authors, a general framework for constructing mixture models for dealing with the high complexity of corrupt / incomplete pairing information of the data with the goal to support a model fitting problem is missing in the literature. Closely related approaches of utilizing mixture models in applied mathematics with lost pairing information are presented for computer vision in order to deal with unknown feature correspondences [Hoegele, 2024a] and for random equations with high combinatorial possibilities for the stochastic parameters [Hoegele, 2024b].
@@
In this paper, we propose to consider the problem of model fitting in a new conclusive way. A general probabilistic framework for fitting models in data based purely on observational probability density functions including errors-in-variables is presented, which has direct relations to well-known standard methods for completely paired data, such as ordinary least squares, Deming regression [Deming, 1964], interval data regression as well as multiple linear regression. This framework will be generalized from supervised to semi-supervised model fitting by including (partially) unpaired data in one common line of stochastic argumentation. The derivations allow for flexibility about i) the number and dimensions of the input / output data, ii) the type of individual error characteristics of each data point with errors-in-variables utilizing general density functions, iii) the type of (linear or nonlinear) models which should be fitted and iv) the pairing information level. In summary, there are two main goals of this paper: A) presenting a general stochastic argumentation framework for model fitting (without ad-hoc loss functions) and B) presenting an extension of model fitting within this framework to partially unpaired data utilizing mixture models.
@@
  • completely paired data sets as the standard case in model fitting (i.e., supervised learning) (Section 2.2).
@@
  • partially unpaired data sets, which lie between the two previous extremes and include different levels of pairing (i.e., semi-supervised learning) (Section 2.4).
@@ 2.2 Model Fit with Completely Paired Data
We introduce disturbances in the input and output data with the notation (which is referred in literature to errors-in-variables [Markovsky and Van Huffel, 2007]):
@@
2.2.1 Example: Fitting a Line and Gaussian Disturbance
In the line of total least squares [Markovsky and Van Huffel, 2007], we are introducing Gaussian disturbances by ${\eta}_{l}\sim\mathcal{N}(0,\,\sigma_{{\eta}}^{2})(s)\;\forall l=1,..,L$ and ${\varepsilon}_{l}\sim\mathcal{N}(0,\,\sigma_{{\varepsilon}}^{2})(s)\;\forall l=1,..,L$. Utilizing the one-dimensional affine model $M(x;\boldsymbol{\alpha})=\alpha_{1}+\alpha_{2}\cdot x$ and inserting it in Equation (4), we get
@@ 2.2.1 Example: Fitting a Line and Gaussian Disturbance
=$\quad$= $\text{argmax}_{\;\boldsymbol{\alpha}}\quad\;\prod\limits_{l=1}^{L}\;\frac{\text{e}^{-\frac{(\alpha_{1}+\alpha_{2}\cdot x_{l}-y_{l})^{2}}{2\,(\alpha_{2}^{2}\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2})}}}{\sqrt{\alpha_{2}^{2}\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2}}}$
=$\quad$= $\text{argmin}_{\;\boldsymbol{\alpha}}\quad\frac{L}{2}\,\ln\left(\alpha_{2}^{2}\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2}\right)+\sum\limits_{l=1}^{L}\;\frac{(\alpha_{1}+\alpha_{2}\cdot x_{l}-y_{l})^{2}}{2\,(\alpha_{2}^{2}\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2})}\;.$ (6)
Since this is a Deming regression type of problem, the solution to this minimization is also closely related to the classical Deming regression. The similarities and differences are presented in Supplement B.
Further extending this argumentation to hyperplanes for $\boldsymbol{x}_{l}\in\mathbb{R}^{k}$ and ${y}_{l}\in\mathbb{R}$, and ${\boldsymbol{\eta}}_{l}\sim\mathcal{N}(\boldsymbol{0},\,\text{diag}(\sigma_{{\eta},1}^{2},..,\sigma_{{\eta},k}^{2}))(\boldsymbol{s})\;\forall l=1,..,L$ and ${\varepsilon}_{l}\sim\mathcal{N}(0,\,\sigma_{{\varepsilon}}^{2})(s)\;\forall l=1,..,L$ for fitting an affine hyperplane model $M(\boldsymbol{x};\boldsymbol{\alpha})=\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{n}$ leads (utilizing Equation (4)) to the general optimization problem
[…]
=$\quad$= $\text{argmax}_{\;\boldsymbol{\alpha}}\quad\prod\limits_{l=1}^{L}\;\frac{\text{e}^{-\frac{\left(\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{l,n}-y_{l}\right)^{2}}{2\,\left(\left(\sum\limits_{n=1}^{k}\alpha_{n+1}^{2}\,\sigma_{{\eta,n}}^{2}\right)+\sigma_{{\varepsilon}}^{2}\right)}}}{\sqrt{\left(\sum\limits_{n=1}^{k}\alpha_{n+1}^{2}\,\sigma_{{\eta,n}}^{2}\right)+\sigma_{{\varepsilon}}^{2}}}$
=$\quad$= $\text{argmin}_{\;\boldsymbol{\alpha}}\quad\frac{L}{2}\,\ln\left(\left(\sum\limits_{n=1}^{k}\alpha_{n+1}^{2}\,\sigma_{{\eta,n}}^{2}\right)+\sigma_{{\varepsilon}}^{2}\right)+\sum\limits_{l=1}^{L}\;\frac{\left(\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{l,n}-y_{l}\right)^{2}}{2\,\left(\left(\sum\limits_{n=1}^{k}\alpha_{n+1}^{2}\sigma_{{\eta,n}}^{2}\right)\,+\sigma_{{\varepsilon}}^{2}\right)}\;.$
@@ 2.2.3 Connection to Interval Data Regression
Interval data is defined as data for which only the borders of an interval in which the true data point lies are observed. Performing model fitting for such data is an active field of research, e.g. for multilinear linear regression models [Lima Neto and De Carvalho, 2008, Souza et al., 2017].
@@ 2.3 Model Fit with Completely Unpaired Data
In this section, we are losing the property of tuples, i.e. observations of type $\boldsymbol{x}_{h}$ for $h=1,..,H$ and $\boldsymbol{y}_{l}$ for $l=1,..,L$ are unpaired. Only a set of $\boldsymbol{x}_{h}$ and a set of $\boldsymbol{y}_{l}$ observations are available. Please note, even the sizes $H$ and $L$ can be different. Although there is theoretical research about the usability of such broken sampling data sets, e.g. [Bai and Hsing, 2005], obviously such data will only lead to very limited regression results if there are no further assumptions about the involved probability densities since we only have marginal distributions. Nonetheless, we want to introduce a formulation by mixture models for this case, which will later be utilized for partially unpaired data directly. We start this argumentation with possible disturbances in input and output data:
@@
Remark: This fit with completely unpaired data is practically useless. This means, there will be broad, probably non-distinct or multiple maxima in this objective function. Still, this argumentation helps in a theoretical perspective since it is applied directly to the partially unpaired data where we have a range of different levels of pairing information.
@@ 2.4 Model Fit with Partially Unpaired Data
This section introduces the argumentation for partially unpaired data, which is a main result of this paper. For this case, we partition the $H$ observations $\boldsymbol{x}_{h}$ and $L$ observations $\boldsymbol{y}_{l}$ into $r=1,..,R$ disjoint independent groups, i.e. observations of group $r$ of type $\boldsymbol{x}_{h,r}$ for $h=1,..,H_{r}$ and $\boldsymbol{y}_{l,r}$ for $l=1,..,L_{r}$ are unpaired with $H_{r}$ and $L_{r}$ representing the number of elements in subgroup $r$. This means, we have a set of $\boldsymbol{x}_{h}$-values and a set of $\boldsymbol{y}_{l}$-values for each subgroup and we have pairing information on the group level. The number of input and output elements in each subgroup $H_{r}$ and $L_{r}$ are not necessarily the same. Again, we allow disturbances in $\boldsymbol{x}_{h}$ and $\boldsymbol{y}_{l}$:
@@
Remark: In the partially unpaired setup, we always work with input / output correspondences, only on a subgroup basis. A completely paired data subset (= supervised data) is represented by subgroups of size one. An additional pure input data subset (= unsupervised) can be approximated by neglecting the information about $\boldsymbol{Y}^{\ast}_{r}$, which corresponds to extremely flat $f_{\boldsymbol{\varepsilon}_{l,r}}$, arriving at the classical definition of semi-supervised. This means, in Equation (11) the first term in the integral $f_{\boldsymbol{\varepsilon}_{l,r}}$ gets essentially constant (independent of the prediction $\boldsymbol{M}(\boldsymbol{s};\boldsymbol{\alpha})$) and, therefore, this part becomes practically noninformative with respect to the optimization on $\boldsymbol{\alpha}$. Thus, as expected, the completely unsupervised part of the data only on the input side can be neglected since it contains no information about the model parameters $\boldsymbol{\alpha}$.
Remark: An interesting point is that the same cannot be said about having unsupervised data on the output side, e.g. neglecting information about $\boldsymbol{X}^{\ast}_{r}$. This time $f_{\boldsymbol{\eta}_{h,r}}$ becomes a flat distribution, essentially leaving $f_{\boldsymbol{\varepsilon}_{l,r}}(\boldsymbol{y}_{l,r}-\boldsymbol{M}(\boldsymbol{s};\boldsymbol{\alpha}))$ in the integral of Equation (11) evaluated for all possible $\boldsymbol{s}$. This time the change of $\boldsymbol{\alpha}$ can have direct influence on the optimization, essentially taking care that the output values $\boldsymbol{y}_{l,r}$ are plausible / possible (i.e. in the probabilistically blurred image of $\boldsymbol{M}(\boldsymbol{s};\boldsymbol{\alpha})$) for a given parameter set $\boldsymbol{\alpha}$. This shows an asymmetry with respect the classical semi-supervised setup [Liang et al., 2007].
2.4.1 Example: Fitting a Line and Gaussian Disturbance